40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Apr 8 22:47:03 2021, shruti, DailyProgress, PSOMA, Cavity locking and temperature scan TemptuneCalc.pdfUntitled.png
    Reply  Wed Apr 14 17:15:23 2021, shruti, DailyProgress, PSOMA, RoC measurement RoCis0_4mConvex.pdf20210412.ziptest_roc.m
       Reply  Thu Apr 15 19:24:35 2021, shruti, DailyProgress, PSOMA, RoC measurement RoC.pdftest_roc.mconcave_beamprofs.zip
Message ID: 2695     Entry time: Thu Apr 8 22:47:03 2021     Reply to this: 2699
Author: shruti 
Type: DailyProgress 
Category: PSOMA 
Subject: Cavity locking and temperature scan 

[Rana, Shruti]

Today Rana pointed out several improvements to the setup including fixing some finicky knob situations. The cavity and setup are now in a good state with monitor and PD connected, although we still have not seen resonance (Will be Attachment 2 soon).

According to the calculations in Attachment 1 a 1V change in the 'TEMP TUNE' input of the ITC-502 temperature (and laser current) controller would correspond to a change of 1 GHz in laser frequency at the current operating point (8 kOhms 'T_ACT' as displayed).

We send in a 1 Hz triangle wave with V_pp=1V into the TEMP TUNE input and make tiny changes to the alignment while monitoring the monitor for flashes and the DC reflected power for a steep drop.

RXA edit:

  • We measured the input coupler transmission to be 2.5% for S-pol, so the cavity Finesse = 2*pi/(1 - T) ~ 250
  • From the overhead picture, you can see that the cavity length is ~15-16 screw holes, so L_{RT} = 16" = 40 cm. So the FSR = c/0.4 ~ 1 GHz.
  • So, assuming the cavity linewidth is then FSR/Finesse = 3 MHz, we want to make sure that our scan velocity is no faster than ~ 1 GHz/s, so that we are able to see the resonances easily on a 1 s oscope trace.
  • The diamter of the beam transmitted through the curved mirror of the cavity is ~3 cm on the monitor.
  • The mode-matching is still far off, since the beam after one round trip is much bigger than the input, but I estimate the mode matching is ~30-50%. Good enough for getting some flashes.
  • It would be good to find a mount for the 1" optics that allows us to cleanly get the 4-ports without clipping.

Other To Do

  • Get a new power adapter for the camera (weird flashes occur when shaking it, confuses the alignment/locking process)
  • Find/purchase a thread adapter to attach longpass filter to Watec camera
  • Eventually organize and safely store all ND filters and other optics in the cabinet
Attachment 1: TemptuneCalc.pdf  1.545 MB  Uploaded Thu Apr 8 23:49:02 2021  | Hide | Hide all
TemptuneCalc.pdf
Attachment 2: Untitled.png  18.535 MB  Uploaded Fri Apr 9 17:56:42 2021  | Hide | Hide all
Untitled.png
ELOG V3.1.3-